# Incommensurate Twisted Bilayer Graphene: emerging quasi-periodicity and stability

### v1

@@ Abstract
We consider a lattice model of Twisted Bilayer Graphene (TBG). The presence of incommensurate angles produces an emerging quasi-periodicity manifesting itself in large momenta Umklapp interactions that almost connect the Dirac points. We rigorously establish the stability of the semimetallic phase via a Renormalization Group analysis combined with number theoretical properties of irrationals, similar to the ones used in Kolmogorov-Arnold-Moser (KAM) theory for the stability of invariant tori. The interlayer hopping is weak and short ranged and the angles are chosen in a large measure set. The result provides a justification, in the above regime, to the effective continuum description of TBG in which large momenta interlayer interactions are neglected.
@@ I Introduction
The discovery that at certain angles Twisted Bilayer Graphene (TBG) develops superconductivity a has generated much interest in such materials both for technological and theoretical reasons b . It was predicted, using continuum models obtained by keeping only the dominant harmonics in the lattice model 1 , 2 ,3 , that at such angles some strongly correlated behaviour should appear, but not of superconducting type. The mechanism behind the superconductivity remains elusive.
Taking the lattice into account breaks several symmetries of the continuum description 1aa1 , 1aa2 leading to effects like the possible shift of Fermi points. More interestingly, for generic angles, excluding a special set 1bb , 1aaa , one has an incommensurate structure; in such a case Bloch band theory does not apply and one has an emergent quasi-periodicity H1 -H5 , with some feature in common with the one in fermions with quasi-periodic potentials P0 -P2 . It is known that electronic quasi-periodic systems have remarkable properties. In 1d they produce a metal-insulator transition Au ,DS ,FS . The interplay with a many body interaction produces peculiar phases with anomalous gaps or many body localization Q0 M11 . Quasi-periodicity has been studied also in Weyl semimetals P4 , W0 ,W1 or in the 2d Ising model Lu , Ca , Ga . It is therefore natural to expect that quasi-periodicity plays an important role in the interacting phases of TBG. Most theoretical analyis are however based on continuum effective description, do not distinguish between commensurate and incommensurate angles, and are based on the assumption that lattice effects preserve the semimetallic phase 1 , 2 , 3 .
[…]
It is expected that the relevant processes in TBG are the ones connecting the Dirac points as closely as possible, that is the terms that minimize the quantity $|G+G^{\prime}+p_{F,i}-p^{\prime}_{F,j}|$ where $p_{F,i}$ $p^{\prime}_{F,j}$ are the Dirac points of the two layers. The approximation at the basis of the effective models 1 , 2 ,3 consists in taking restricting the interaction to only the terms $G=G^{\prime}=0$ or $G=b_{1},G^{\prime}=-b^{\prime}_{1}$ or $G=b_{2},G^{\prime}=-b^{\prime}_{2}$ and taking the continuum limit, based on the fact that larger values of $G$ or $G^{\prime}$ are exponentially depressed 111 . However in the incommensurate case, Umklapp terms with very large values of $G,G^{\prime}$ make $|G+G^{\prime}+p_{F,i}-p^{\prime}_{F,j}|$ arbitrarely small, producing almost relevant processes which can destroy the semimetallic behaviour. In the 1d Aubry-André model the processes that produce small values for $2\varepsilon p_{F}+2l\pi+2\pi\omega m$, $\varepsilon=0,\pm 1$ are indeed the ones producing the insulating behaviour at large coupling, while at weak coupling the metallic regime persists. Similarly the persistence or not of the semimetallic regime in TBG depends on the relevance or irrelevance of the terms involving large $G,G^{\prime}$ that almost connect the Dirac points. This fact cannot be understood only on the basis of perturbative arguments; it is indeed a non perturbative phenomenon which can be established only by the convergence or divergence of the whole series expansion. Despite the similarity of quasi-periodic potentials and incommensurate TBG, there are crucial differences like the higher dimensionality of TBG and the fact that the frequencies are not independent parameters but are functions of a single parameter, the angle between the layers, and this produces rather different small divisors.
The aim of this paper is to investigate when the quasi-metallic phase is stable against the large momentum processes in the incommensurate case. The analysis is based on Renormalization Group methods combined with number theoretical properties of irrationals, similar to the ones used in Kolmogorov-Arnold-Moser (KAM) theory for the stability of invariant tori. Due to the difficulty of getting information on the single particle spectrum, we analyze the behavior of the Euclidean correlations, which provide information on the spectrum close to the Fermi points. Such methods are robust enough to be extended to many-body systems, as it was done for the interacting Aubry-André model M11 or in Weyl semimetals W1 . Our main result is the proof of the stability of the semimetallic phase in a large measure set of angles in the incommensurate case.
@@
The paper is organized in the following way. In Section II the lattice model of TBG is presented. In Section III a perturbative expansion for the correlations is derived. In Section IV the emerging quasi-periodicity and the small divisor problem is described, together with the required (number theoretical) Diophantine conditions. Section V contains a statement of the main result and in Section VI the Renormalization Group derivation is presented. The Appendices detail the more technical aspects of the analysis.
@@ II Incommensurate TBG
We consider the lattice TBG model introduced in 1 , 2 . We focus on this model for the sake of definiteness but our methods could be applied more generally. We consider two graphene layers rotated with respect to one another by an angle $\theta$ around a point $\xi=(0,1/2)$ (that is, the point between an a and b atom, chosen so that the twisted model preserves the $C_{2}T$ symmetry in Appendix D.1). The Hamiltonian of the system will be written as
@@
where $H_{1}$ and $H_{2}$ are hopping Hamiltonians within the layers 1 and 2 respectively and $V$ is an interlayer hopping term. The first graphene layer is defined on the lattice $\mathcal{L}_{1}:=\{n_{1}A_{1}+n_{2}A_{2},\ n_{1},n_{2}\in\mathbb{Z}\}$ with $A_{1}={1\over 2}(3,\sqrt{3}),\quad A_{2}={1\over 2}(3,-\sqrt{3})$. We introduce the nearest-neighbor vectors: $\delta_{1}=(1,0)$, $\delta_{2}={1\over 2}(-1,\sqrt{3})$, $\delta_{2}={1\over 2}(-1,-\sqrt{3})$. We will write the Hamiltonian in second quantized form: for $x\in\mathcal{L}_{1}$, we introduce annihilation operators $c_{1,x,a}$ and $c_{1,x,b}$ corresponding respectively to annihilating a fermion located at $x$ and $x+\delta_{1}$. The nearest neighbor hopping Hamiltonian is
@@
with $|\widehat{\mathcal{L}}_{1}|=8\pi^{2}/3\sqrt{3}$, and $\widehat{\mathcal{L}}_{1}:=\mathbb{R}^{2}/(b_{1}\mathbb{Z}+b_{2}\mathbb{Z})$ in which
$b_{1}={\textstyle{2\pi\over 3}}(1,\sqrt{3}),\quad b_{2}={\textstyle{2\pi\over 3}}(1,-\sqrt{3}).$ (4)
In Fourier space,
@@
with $\Omega(k_{x},k_{y}):=1+2e^{-i\frac{3}{2}k_{x}}\cos({\textstyle\frac{\sqrt{3}}{2}k_{y}})$.
@@
In the absence of interlayer coupling the two graphene layers are decoupled; the single particle spectrum for layer 1 is $\pm|\Omega(k)|$ and the Fermi points are given by the relation $\Omega(p_{F,1}^{\pm})=0$ with
$p_{F,1}^{\pm}={2\pi\over 3}(1,\pm{\textstyle{1\over\sqrt{3}}})$ (10)
for momenta close to such points one has $|\Omega(k)|\sim{3\over 2}t|k-p_{F,1}^{\pm}|$, that is the dispersion relation is almost linear (relativistic) up to quadratic corrections, forming approximate Dirac cones. In the same way the dispersion relation for layer 2 is $\pm|\Omega(R^{T}k)|$; the Fermi points are $\Omega(R^{T}p_{F,2}^{\pm})=0$ with $p_{F,2}^{\pm}=R(p_{F,1}^{\pm})$ and $|\Omega(R^{T}k)|\sim{3\over 2}t|k-p_{F,2}^{\pm}|$. We are interested in understanding how these four Dirac cones are modified in the presence of the interlayer hopping.
@@
We restrict the interlayer term to hoppings between atoms of type a to atoms of type a and type b to b for technical reasons. This is so that the “Inversion” symmetry in Appendix D.1 is satisfied. We could just as easily consider a model where the interlayer hopping occurs only between atoms of type a to atoms of type b. We could relax this restriction and allow for all possible hoppings, but we would then need to add extra counterterms (see (15)) to the model. For the sake of simplicity, we avoid this and only consider these interlayer hoppings.
Note that, whereas the Fourier transform for $c$ is defined on $\widehat{\mathcal{L}}_{i}$, the Fourier transform of $\varsigma$ is defined on all $\mathbb{R}^{2}$. We write $V$ in Fourier space: we get, see App. A
@@
$\left.+\sum_{m\in\mathbb{Z}^{2}}\int_{\widehat{\mathcal{L}}_{2}}dk\ \tau_{m,\alpha}^{(2)}(k+mb^{\prime})\widehat{c}^{\dagger}_{2,k,\alpha}\widehat{c}_{1,k+mb^{\prime},\alpha}\right)$
@@
The position of the Dirac points are in general modified (renormalized) by the interlayer hopping. It is conventient to fix the values of the renormalized Dirac points by properly choosing the bare ones. This can be achieved by replacing $\Omega(k)$ with $\Omega(k)+\nu_{i,\omega}$ close to each Dirac points, that is adding a counterterm has the form
@@
where $\chi_{\omega,i}(k)$ is a smooth compactly supported function that is non vanishing for $k-p_{F,i}^{\omega}||_{i}\leq 1/\gamma$, for some $\gamma>1$, in which $||.||_{i}$ is the norm on the torus $\widehat{\mathcal{L}}_{i}$.
Our main result concerns the two-point Schwinger function, which we define as follows. We first introduce a Euclidean time component: given an inverse temperature $\beta>0$, we define for $x_{0}\in[0,\beta)$,
@@
which is defined for $(k_{0},k)\in\widehat{\Lambda}_{j}:=\frac{2\pi}{\beta}\mathbb{Z}\times\widehat{\mathcal{L}}_{j}$.
Now, given $j,j^{\prime}\in\{1,2\}$, $\mathbf{k}=(k_{0},k)\in\Lambda_{j}$, the two-point Schwinger function is defined as the $2\times 2$ matrix $S_{j,j^{\prime}}(\mathbf{k})$ whose components are indexed by $\alpha,\alpha^{\prime}\in\{a,b\}$:
@@
To compute the Schwinger functions, we will use the Grassmann integral formalism. To do so, we add an imaginary time component to all position vectors: we define $\mathbf{A}_{i}:=(0,A_{i})$. Given $\mathbf{x}=(x_{0},x)\in\Lambda_{j}\equiv[0,\beta)\times\mathcal{L}_{j}$, we introduce Grassmann variables $\psi^{\pm}_{\mathbf{x},a,j},\psi^{\pm}_{\mathbf{x},b,j}$ and their Fourier transforms
@@
with $|\widehat{\Lambda}_{j}|=8\pi^{2}/3\sqrt{3}$, $\bm{\xi}=(0,\xi)$ and $\widehat{\Lambda}_{j}=\frac{2\pi}{\beta}\mathbb{Z}\times\widehat{\mathcal{L}}_{j}$. The Schwinger fiunctions are given by
$(S_{j,j^{\prime}}({\bf x},{\bf y}))_{\alpha,\alpha^{\prime}}:={\int P(d\psi)\ e^{-\beta(V(\psi)+M(\psi))}\psi^{-}_{j,{\bf x},\alpha}\psi^{+}_{j^{\prime},{\bf y}\,\alpha^{\prime}}\over\int P(d\psi)\ e^{-\beta(V(\psi)+M(\psi))}}$ (21)
@@
$\widehat{g}_{1}({\bf k}+p_{F}^{\pm})\sim\begin{pmatrix}-ik_{0}&-v_{F}(-ik_{1}\pm k_{2})\\ -v_{F}(ik_{1}\pm k_{2})&-ik_{0}\end{pmatrix}^{-1}.$ (23)
@@
with $\widehat{g}_{2}({\bf k})=\begin{pmatrix}-ik_{0}&\Omega(R^{T}k)\\ \Omega^{*}(R^{T}k)&-ik_{0}\end{pmatrix}^{-1}\equiv\widehat{g}_{1}(R^{T}\mathbf{k})$ which is singular at $p_{F,2}^{\pm}:=Rp_{F,1}^{\pm}$ and periodic in $k$ with period $b^{\prime}_{1},b^{\prime}_{2}$. The interaction and counterterm are rewritten formally in terms of Grassmann variables:

### v2

@@ Abstract
We consider a lattice model of twisted bilayer graphene (TBG) for incommensurate twist angles, focusing on the role of large-momentum-transfer Umklapp terms. These terms, which nearly connect the Fermi points of different layers, are typically neglected in effective continuum descriptions but could, in principle, destroy the Dirac cones; they are indeed closely analogous to those appearing in fermions within quasi-periodic potentials, where they play a crucial role. We prove that, for small but finite interlayer coupling, the semimetallic phase is stable provided the angles belong to a fractal set of large measure (which decreases with the hopping strength) characterized by a number-theoretic Diophantine condition. In particular, this set excludes the (zero measure) commensurate angles. Our method combines a Renormalization Group (RG) analysis of the imaginary-time, zero-temperature Green’s functions, with number theoretic properties, and it is similar to the technique used in the Lindstedt series approach to Kolmogorov-Arnold-Moser (KAM) theory. The convergence of the resulting series allows us to rule out non-perturbative effects. The result provides a partial justification of the effective continuum description of TBG in which such large-momentum interlayer hopping processes are neglected.
@@ I Introduction
Twisted Bilayer Graphene (TBG) consists of two graphene layers, stacked and twisted by an angle $\theta$. The electronic properties of graphene are accurately captured by a tight-binding model of electrons hopping from one site to a neighboring one of a honeycomb lattice; the corresponding energy bands can be explicitly computed and intersect at two Dirac points, close to which the effective dispersion relation is approximately conical. In the case of half-filling, low-energy excitations have an effective continuum description in terms of two-dimensional massless Dirac fermions and the system behaves as a Dirac semimetal.
Stacking and twisting two honeycomb lattices produces an intricate Moiré pattern [1] and twisted bilayer graphene behaves very differently from the single layer, with a strong sensitivity to the twist angle, ranging from non-conductive to superconductive [2], see e.g. [3],[4],[5]. From the theoretical side, in contrast to the monolayer case, an exact computation of the dispersion relation is impossible and the validity of effective models is less clear.
The model for TBG we will consider consists of two layers of graphene with a quadratic interlayer coupling $\lambda$ smaller than the intralayer hopping (chosen equal to $1$). We denote the twist angle by $\theta$. One has to distinguish between commensurate angles, such that the TBG is periodic and which are a discrete zero measure set indexed by two integers [6],[7],[8]; and incommensurate angles where periodicity is lost and Bloch theory cannot be applied [9].
The Dirac points of layer $1$ are $K_{\pm}={2\pi\over 3}(1,\pm{\textstyle{1\over\sqrt{3}}})$ and the reciprocal lattice vectors are $b_{1},b_{2}$; correspondingly the Dirac points of layer 2 are $K^{{}^{\prime}}_{\pm}=R(\theta)K_{\pm}$ in which $R(\theta)$ is the rotation matrix of angle $\theta$, and the reciprocal lattice vectors are $b^{{}^{\prime}}_{i}=R(\theta)b_{i}$, $i=1,2$. The interlayer hopping with coupling $\lambda$ produces a sum of terms with an exchange of momenta given by multiples of the reciprocal lattice vectors; if $k_{a},k_{b}$ are the momenta of two fermions, the conservation rule is given by
$k_{a}-k_{b}+G+G^{\prime}=0$ (1)
with $G=l_{1}b_{1}+l_{2}b_{2}=lb$, $G^{\prime}=m_{1}b^{\prime}_{1}+m_{2}b_{2}^{\prime}=mb^{\prime}$, $l=(l_{1},l_{2})$, $m\equiv(m_{1},m_{2})$ with $l_{i},m_{i}$ integers. The processes involving non-vanishing $G+G^{\prime}$ are called Umklapp terms. The interlayer hopping decreases with large exchange of momentum.
The processes connecting the Dirac points are of particular importance, that is the processes such that $\Delta=0$ where
$\Delta=|G+G^{\prime}+K_{\sigma}^{i}-K_{\sigma^{\prime}}^{j}|\equiv|M+K_{\sigma}^{i}-K_{\sigma^{\prime}}^{j}|$ (2)
with $K_{\pm}^{1}=K^{\pm}$, $K_{\pm}^{2}=K^{{}^{\prime}}_{\pm}$ and $M=G+G^{\prime}$. Among such terms, one can distinguish between the ones connecting the same Dirac point (that is $i=j,\sigma=\sigma^{\prime}$) from those connecting different Dirac points. The form of the former is determined by the lattice symmetries [10, 11] and their effect is to renormalize the parameters of the single layer Hamiltonian, possibly enlarged to include new terms allowed by the breaking of symmetries produced by the interlayer hopping. The latter, that is the ones involving different Dirac points, are present only for commensurate angles and are expected to be relevant and able to destroy the Dirac cones [12]. One can distinguish between the intralayer terms ($i=j,\sigma\not=\sigma^{\prime}$) and interlayer terms ($i\not=j,\sigma\not=\sigma^{\prime}$); the first case corresponds to the opening of a gap and the other to quadratic bands. The gap is expected to decrease with the momentum transferred due to the decay in momentum of the hopping. Due to periodicity there is only a finite set of exchanged Umklapp momenta, see Figure 1.
Figure 1: Left: The Dirac points in the commensurate case $\cos\theta=37/38$. Right: The Dirac points in the approximately incommensurate case $\theta=0.16\pi$. The Dirac points $K_{\pm}$ of one layer are in blue (color online), and those of the other layer $K_{\pm}^{\prime}+G+G^{\prime}$ are in red. In the incommensurate case, $K^{\prime}_{\pm}+G+G^{\prime}$ get arbitrarily close to $K_{\pm}$.
Therefore, in general, in the commensurate case the terms that do not connect the Dirac points exactly have a large mismatch and are expected to be irrelevant.
Also the terms almost connecting the Dirac points can produce an important modification of the physical properties with respect to graphene. Terms almost connecting the Dirac points can be of two kinds. A first possibility is for angles $\theta$ close to zero, where one can see that $\Delta$ can be $O(\theta)$; they are present both in the commensurate or incommensurate case. An effective model was proposed in [13],[14], [15] keeping only a few terms ($G=G^{\prime}=0$, or $G=b_{1},G^{\prime}=-b^{\prime}_{1}$ or $G=b_{2},G^{\prime}=-b^{\prime}_{2}$) corresponding to $\Delta=O(\theta)$ and passing to the continuum. It was found that the effective hopping is greatly enlarged: it becomes of order $O(\lambda/\theta)$; therefore the conical dispersion relation can persist but with a strongly decreased velocity; for certain ”magic” angles the velocity even vanishes and, remarkably, superconductivity has been found [2]. Effective models of TBG in which not only the velocity vanishes but the bands are completely flat have been also proposed [16] and mathematical derivations of such effective models in suitable limits have been provided [17].
The other possibility is present only in the incommensurate case and involves large momentum transfer. Indeed in this case there are infinitely many Umklapp terms almost connecting the Dirac points, that is such that the distance between $K_{\pm}$ and $K_{\pm}^{\prime}+G^{\prime}+G$ can be arbitrarily small, that is $\Delta$ can be arbitrarily small, see figure 1. These terms, which nearly connect the Fermi points of different layers, are typically neglected in effective continuum descriptions but could, in principle, destroy the Dirac cones; they are indeed closely analogous to those appearing in the case of fermions in quasi-periodic potentials, where they play a crucial role. This can be easily realized comparing TBG for incommensurate angles with fermions on a chain with step $1$ and with an Aubry quasi-periodic potential $\cos 2\pi\omega x$ [18], whose single partcle properties are described by the Harper equation. In that case the conservation of momentum reads $k_{a}-k_{b}+2l\pi+2\pi\omega m=0$ and the mismatch between Fermi points is
$\Delta=|2\varepsilon K+2l\pi+2\pi\omega m|\quad\varepsilon=0,\pm 1$ (3)
if $K$ is the Fermi momentum. If $\omega/2\pi$ is rational either the above quantity is zero or is bounded while if it is irrational there are almost relevant processes that almost connect the Fermi points. The quasi-periodicity acts as a disorder and in the strong coupling regime there is an insulating localized behaviour while in the weak coupling a metallic behaviour is found, as proved in [19], [20], [21]. The spectrum has remarkable fractal properties, which were observed in experiments with magnetic fields [22].
The prototypical problem involving quasi-periodicity is the Kolmogorov-Arnold-Moser (KAM) theorem for the persistence of Hamiltonian invariant tori under perturbation, and, in fact, results on the Aubry potential use KAM methods. The Renormalization Group can be used to prove the convergence of the KAM Lindstedt series [23] or to compute the Green’s function of fermions in a quasi periodic potential [24]. It is also known that the interplay of many body interaction and quasi-periodic potential leads to interesting phenomena, studied in $1d$ fermions in the delocalized regime in [25], [26] and in the localized one in [27], [28]. Quasi-periodic potentials have also been investigated in interacting Weyl semimetals [29], [30]. Note also that the large-momentum Umklapp terms in incommensurate TBG are also present when considering a hopping with a finite number of harmonics; they would be generated at higher order in perturbation theory. This fact is not captured by the effective models, as the lattice is replaced by the continuum and Umklapp terms are therefore absent. These large-momentum Umklapp terms are depressed by the decay of the hopping in momentum space, but there is no guarantee that this makes them irrelevant; in the case of the Aubry potential they indeed produce an Anderson localized phase.
In order to evaluate the effect of the Umklapp large-momentum terms on the physical properties of the system we need, as is the case in KAM theory or for quasi-periodic systems, to use number theoretic properties of irrationals. Such properties are typically established in the context of measure theory. The typical number theoretic condition which is assumed in the case of problems involving quasi-periodicity is a Diophantine condition which consists in restricting ourselves to values of $\theta$ for which, for any $\omega,\omega^{\prime},i,j$,
$|K^{i}_{\omega}-K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|\geq\frac{C_{0}}{|y|^{\tau}},$ (4)
where $y$ is either $y=l$ or $y=m$ and $y\neq 0$; this relates the size of the momentum $G,G^{\prime}$ with the precision with which the Dirac points are connected. A similar Diophantine condition is assumed in the case of Aubry-André systems, where the condition is on the frequency $\omega$, and in KAM theory where the condition is on the frequency of the tori. $C_{0}$ is called the Diophantine constant and it plays a crucial role.
We consider a lattice model of TBG and we perform an exact Renormalization Group analysis of the 2-point zero temperature imaginary time Green function in order to find conditions for the persistence of the Dirac cones and the stability of the semimetallic phase. Such zero temperature correlations provide information on the ground state and in particular on the persistence or not of the semimetallic behaviour. Note that, as we are not considering a many-body interaction, information on the cones could also be obtained, in principle, by studying the corresponding one-particle Schrödinger operator. We find however more suitable the functional integral formalism, considering the hopping as a perturbation and expanding the correlation as a convergent series.
The stability is a non-perturbative phenomenon which cannot be understood on the basis of lowest order computations but it depends on the convergence or divergence of the whole series. This is a common feature in problems involving incommensurability; the KAM series are convergent and correspond to stable motions while the Birkhoff series are diverging and correspond to chaos but both have a small divisor problem.
We choose an interlayer hopping with coupling $\lambda$ that decays exponentially for large momenta and satisfies suitable lattice symmetries so that the terms connecting the same Dirac points just renormalize the single layer parameters except for a modification of the location of Dirac points. Such a choice of symmetries is done for definiteness and it could be avoided by including other terms (counterterms) in the single layer Hamiltonian that are compatible with the symmetries, while the fast decay of the hopping in momentum space seems essential.
Our main result is the following:
In a lattice model of twisted bilayer graphene with an angle $\theta$ verifying the Diophantine condition (4) for $|\lambda|\leq\varepsilon_{0}(C_{0})$ with $\varepsilon_{0}(C_{0})$ proportional to $C_{0}^{\beta}$ for some positive power $\beta$, the Dirac cones persist and the Dirac semimetallic phase is stable.
This result ensures that the large momentum Umklapp terms, even if they almost connect the Dirac points, are not able to destroy the semimetallic phase at weak coupling. The proof is based on a Renormalization Group analysis for the zero temperature imaginary time Green functions and the convergence of the series allows to exclude non-perturbative effects.
The stability of the cones is established with a smallness condition which depends on a power of the Diophantine constant $C_{0}$; given an angle one can in principle compute it but it is expected to depend wildly on the exact value of $\theta$. It is therefore convenient to give a generic statement for angles in a certain set, as it is done for disordered systems where the properties hold for generic (not all) configurations of disorder. Given an angle in a certain interval of size $2\,\delta\theta$ around $\theta$, we can prove that the angles verifying (4) have almost a full measure: the complement of the set has a measure that is bounded by $C_{\theta,\delta\theta}C_{0}^{\alpha}(2\,\delta\theta)$, with some positive power $\alpha$ and a constant $C_{\theta,\delta\theta}$ (that depends on derivatives of $K^{\pm}_{i},b_{j}$); the explicit value of the power $\alpha$ and the form of $C_{\theta,\delta\theta}$ are written explicitly below. One can check that $C_{\theta,\delta\theta}$ is finite apart from a finite set of special, pathological angles (there are 20 such pathological angles, which arise from technical issues). Therefore choosing a hopping $\lambda$ fixes an admissible value for $C_{0}$, hence choosing, say, $\lambda=\varepsilon_{0}(C_{0})/2$ the measure of the angles for which the Dirac cones persist has a complement that is bounded by $C_{\theta,\delta\theta}C_{0}^{\alpha}(2\,\delta\theta)$. In addition $\delta\theta$ also has to be chosen small with respect to $C_{0}$ but this is probably just a technical assumption which could be removed by a lengthier analysis. The set for which stability is ensured is of course rather complex and has a fractal nature and its relative measure decreases with the hopping. The situation is therefore entirely analogous to the breaking of invariant tori in KAM theory, see e.g. [31], [32]. Note that the excluded set contains the commensurate angles, but experimentally to see an instability can be difficult as the possible gap could be very small if the corresponding Umklapp momentum is very large. Another equivalent way to state the result is that for any angle (except the pathological ones) one can find an interval of angles near it such that, with probability $1$, there is a non-zero value of the hopping for which the semimetallic behaviour persists.
Our result therefore ensures that for weak hopping for most (in the sense specified above) angles the semimetallic behaviour persists even in presence of the large momentum Umklapp terms which almost connect the Dirac points. The result provides a partial justification of the effective continuum description of TBG in which terms are neglected. It would be of course a very interesting problem to consider hopping outside the regime we are considering to see if there can be some transition to a different phase due to such terms, as it happens for fermions in a 1D quasi-periodic Aubry potential; or to detect in the weak coupling regime some sign of fractality following from the above number theoretical considerations.
@@
The paper is organized in the following way. In Section II the lattice model of TBG is presented. In Section III a perturbative expansion for the correlations is derived. The main result is stated in Section IV. In Section V the emerging quasi-periodicity and the small divisor problem is described, together with the required (number theoretic) Diophantine conditions. In Section VI the Renormalization Group derivation is presented. The Appendices detail the more technical aspects of the analysis.
@@ II Incommensurate TBG
We consider the lattice TBG model introduced in [13, 14]. We focus on this model for the sake of definiteness but our methods could be applied more generally. We consider two graphene layers rotated with respect to one another by an angle $\theta$ around a point $\xi=(0,1/2)$ (that is, the point between an a and b atom, chosen so that the twisted model preserves the $C_{2}T$ symmetry in Appendix D.1). The Hamiltonian of the system will be written as
@@
where $H_{1}$ and $H_{2}$ are hopping Hamiltonians within the layers 1 and 2 respectively and $V$ is an interlayer hopping term. The first graphene layer is defined on the lattice $\mathcal{L}_{1}:=\{n_{1}A_{1}+n_{2}A_{2},\ n_{1},n_{2}\in\mathbb{Z}\}$ with $A_{1}={1\over 2}(3,\sqrt{3}),\quad A_{2}={1\over 2}(3,-\sqrt{3})$. We introduce the nearest-neighbor vectors: $\delta_{1}=(1,0)$, $\delta_{2}={1\over 2}(-1,\sqrt{3})$, $\delta_{2}={1\over 2}(-1,-\sqrt{3})$. We will write the Hamiltonian in second quantized form: for $x\in\mathcal{L}_{1}$, we introduce annihilation operators $c_{1,x,a}$ and $c_{1,x,b}$ corresponding respectively to annihilating a fermion located at $x$ and $x+\delta_{1}$. The nearest neighbor hopping Hamiltonian is ($t=1$ from now on)
@@
with $|\widehat{\mathcal{L}}_{1}|=8\pi^{2}/3\sqrt{3}$, and $\widehat{\mathcal{L}}_{1}:=\mathbb{R}^{2}/(b_{1}\mathbb{Z}+b_{2}\mathbb{Z})$ in which $b_{1}={\textstyle{2\pi\over 3}}(1,\sqrt{3}),\quad b_{2}={\textstyle{2\pi\over 3}}(1,-\sqrt{3})$ In Fourier space,
@@
with $\Omega(k_{x},k_{y}):=1+2e^{-i\frac{3}{2}k_{x}}\cos({\textstyle\frac{\sqrt{3}}{2}k_{y}})$. Note that, for small $\theta$, $|K_{\pm}-K^{{}^{\prime}}_{\pm}|=O(\theta)$. .
@@
In the absence of interlayer coupling the two graphene layers are decoupled; the single particle spectrum for layer 1 is $\pm|\Omega(k)|$ and the Fermi points are given by the relation $\Omega(K^{1}_{\pm})=0$ with
$K^{1}_{\pm}={2\pi\over 3}(1,\pm{\textstyle{1\over\sqrt{3}}})$ (13)
for momenta close to such points one has $|\Omega(k)|\sim{3\over 2}t|k-K^{1}_{\pm}|$, that is the dispersion relation is almost linear (relativistic) up to quadratic corrections, forming approximate Dirac cones. In the same way the dispersion relation for layer 2 is $\pm|\Omega(R^{T}k)|$; the Fermi points are $\Omega(R^{T}K^{2}_{\pm})=0$ with $K^{2}_{\pm}=R(K^{1}_{\pm})$ and $|\Omega(R^{T}k)|\sim{3\over 2}t|k-K^{2}_{\pm}|$. We are interested in understanding how these four Dirac cones are modified in the presence of the interlayer hopping.
@@
We restrict the interlayer term to hoppings between atoms of type a to atoms of type a and type b to b so that the “Inversion” symmetry in Appendix D.1 is satisfied. Note that, whereas the Fourier transform for $c$ is defined on $\widehat{\mathcal{L}}_{i}$, the Fourier transform of $\varsigma$ is defined on all $\mathbb{R}^{2}$. We write $V$ in Fourier space: we get, see App. A
@@
$\left.+\sum_{m\in\mathbb{Z}^{2}}\int_{\widehat{\mathcal{L}}_{2}}dk\ \tau_{m,\alpha}^{(2)}(k+mb^{\prime})\widehat{c}^{\dagger}_{2,k,\alpha}\widehat{c}_{1,k+mb^{\prime},\alpha}\right)$ (16)
@@
The position of the Dirac points are in general modified (renormalized) by the interlayer hopping. It is convenient to fix the values of the renormalized Dirac points by properly choosing the bare ones. This can be achieved by replacing $\Omega(k)$ with $\Omega(k)+\nu_{i,\omega}$ close to each Dirac points, that is adding a counterterm has the form
@@
where $\chi_{\omega,i}(k)$ is a smooth compactly supported function that is non vanishing for $k-K^{i}_{\omega}||_{i}\leq 1/\gamma$, for some $\gamma>1$, in which $||.||_{i}$ is the norm on the torus $\widehat{\mathcal{L}}_{i}$.
Our main result concerns the Matsubara imaginary time Green functions, which we define as follows. We first introduce a Euclidean time component: given an inverse temperature $\beta>0$, we define for $x_{0}\in[0,\beta)$,
@@
which is defined for $(k_{0},k)\in\widehat{\Lambda}_{j}:=\frac{2\pi}{\beta}(\mathbb{Z}+1/2)\times\widehat{\mathcal{L}}_{j}$.
Now, given $j,j^{\prime}\in\{1,2\}$, $\mathbf{k}=(k_{0},k)\in\Lambda_{j}$, the 2-point function is defined as the $2\times 2$ matrix $\widehat{S}_{j,j^{\prime}}(\mathbf{k})$ whose components are indexed by $\alpha,\alpha^{\prime}\in\{a,b\}$:
@@
To compute the 2-point functions we will use the Grassmann integral formalism. To do so, we add an imaginary time component to all position vectors: we define $\mathbf{A}_{i}:=(0,A_{i})$. Given $\mathbf{x}=(x_{0},x)\in\Lambda_{j}\equiv[0,\beta)\times\mathcal{L}_{j}$, we introduce Grassmann variables $\psi^{\pm}_{\mathbf{x},a,j},\psi^{\pm}_{\mathbf{x},b,j}$ and their Fourier transforms
@@
with $|\widehat{\Lambda}_{j}|=8\pi^{2}/3\sqrt{3}$, $\bm{\xi}=(0,\xi)$. The 2-point function can be written as
$(S_{j,j^{\prime}}({\bf x},{\bf y}))_{\alpha,\alpha^{\prime}}:={\int P(d\psi)\ e^{-V(\psi)-M(\psi)}\psi^{-}_{j,{\bf x},\alpha}\psi^{+}_{j^{\prime},{\bf y}\,\alpha^{\prime}}\over\int P(d\psi)\ e^{-V(\psi)-M(\psi)}}$ (25)
@@
$\widehat{g}_{1}({\bf k}+K^{\pm})\sim\begin{pmatrix}-ik_{0}&-v_{F}(-ik_{1}\pm k_{2})\\ -v_{F}(ik_{1}\pm k_{2})&-ik_{0}\end{pmatrix}^{-1}.$ (27)
@@
with $\widehat{g}_{2}({\bf k})=\begin{pmatrix}-ik_{0}&\Omega(R^{T}k)\\ \Omega^{*}(R^{T}k)&-ik_{0}\end{pmatrix}^{-1}\equiv\widehat{g}_{1}(R^{T}\mathbf{k})$ which is singular at $K^{2}_{\pm}:=RK^{1}_{\pm}$ and periodic in $k$ with period $b^{\prime}_{1},b^{\prime}_{2}$. The interaction and counterterm are rewritten formally in terms of Grassmann variables: